### v3

# Simple and Optimal Sublinear Algorithms for Mean Estimation

Abstract
We study the sublinear multivariate mean estimation problem in $d$-dimensional Euclidean space. Specifically, we aim to find the mean $\mu$ of a ground point set $A$, which minimizes the sum of squared Euclidean distances of the points in $A$ to $\mu$. We first show that a multiplicative $(1+\varepsilon)$ approximation to $\mu$ can be found with probability $1-\delta$ using $O(\varepsilon^{-1}\log\delta^{-1})$ many independent uniform random samples, and provide a matching lower bound. Furthermore, we give two sublinear time algorithms with optimal sample complexity for extracting a suitable approximate mean:
  1. 1.
    A gradient descent approach running in time $O((\varepsilon^{-1}+\log\log\delta^{-1})\cdot\log\delta^{-1}\cdot d)$. It optimizes the geometric median objective while being significantly faster for our specific setting than all other known algorithms for this problem.
  2. 2.
    An order statistics and clustering approach running in time $O\left((\varepsilon^{-1}+\log^{\gamma}\delta^{-1})\cdot\log\delta^{-1}\cdot d\right)$ for any constant $\gamma>0$.
Throughout our analysis, we also generalize the familiar median-of-means estimator to the multivariate case, showing that the geometric median-of-means estimator achieves an optimal sample complexity for estimating $\mu$, which may be of independent interest.
Machine Learning, ICML
[…]
1 Introduction
An extremely simple algorithmic paradigm is to sample a subset of the input independently and uniformly at random and solve a problem on the sample. Such algorithms are called sublinear algorithms and form a staple of large data analysis. The most desirable property of a sublinear algorithm is that the query of sample complexity is as small as possible. For most problems, it is rare for a sublinear algorithm with only constant queries to produce multiplicative guarantees. Remarkably, finding the multivariate mean of a point set $A$ in Euclidean space is one of the cases where not only can we obtain a multiplicative error with constant queries, but can do so with arbitrary precision. Specifically, we are interested in the objective
$\min_{c\in\mathbb{R}^{d}}\sum_{p\in A}\|p-c\|^{2},$
where $\|x\|=\sqrt{\sum_{i=1}^{d}x_{i}^{2}}$ denotes the Euclidean norm. It is well known that the mean of the point set $A$ minimizes this expression. We say that $c$ is a $(1+\varepsilon)$-approximate mean, if the cost of $c$ is within a $(1+\varepsilon)$ factor of the cost of the optimal mean.
The exact sample complexity for obtaining approximate means is still an open problem. Inaba et al. (1994) showed that $\varepsilon^{-1}$ uniform samples are necessary and sufficient for the empirical mean to be a $(1+\varepsilon)$-approximate mean on expectation. To extend this to a high probability guarantee, that is, a success probability of $1-\delta$, $O(\varepsilon^{-1}\delta^{-1})$ samples are sufficient. Unfortunately, this bound is tight when using the empirical mean [^1] We give a proof of this in Section 5.. Cohen-Addad et al. (2021) improved the dependence on the probability of failure by achieving a sample complexity of $\tilde{O}(\varepsilon^{-5}\log^{2}\delta^{-1})$, at the cost of a significantly worse dependency on $\varepsilon$. In the same paper, the authors also gave a lower bound of $\Omega(\varepsilon^{-1})$ for any sublinear algorithm succeeding with constant probability, showing that the result by Inaba et al. (1994) is optimal in the constant success probability regime. Subsequently, (Musco et al., 2022) presented an algorithm with $O\left(\varepsilon^{-1}d(\log^{2}d\cdot\log(d\varepsilon^{-1})+\log\delta^{-1})\log\delta^{-1}\right)$ queries, which combines the tight dependency on $\varepsilon$ from (Inaba et al., 1994) with the improved dependency on $\delta$ from Cohen-Addad et al. (2021), at the cost of a dependency on the dimension. Recently, Woodruff & Yasuda (2024) achieved almost best of all worlds trade-offs with a sample complexity of $O(\varepsilon^{-1}(\log\varepsilon^{-1}+\log\delta^{-1})\log\delta^{-1})$.
1.1 Our Contributions
Our main contributions are two algorithms that estimate the mean of a point set and that have optimal sample complexity. In addition, the running times are almost linear in the sample complexity and thus nearly optimal. In the first step of either algorithm, see also Algorithm 1, we draw $O(\varepsilon^{-1}\log\delta^{-1})$ points uniformly at random. We then split the sample into $O(\log\delta^{-1})$ equally sized subsamples of size $O(\varepsilon^{-1})$, compute the empirical mean for each subsample, and then run an aggregation procedure, i.e., a subroutine $\textsc{Aggregate}(P)$ that combines the sample set produced to output an estimate of the mean.
Algorithm 1 MeanEstimate$(\varepsilon,\delta)$
  for $i=1,\ldots,b\log\delta^{-1}$ do
     Take a uniform sample $S_{i}$ with $|S_{i}|=a\varepsilon^{-1}$ points
     Compute the sample mean $\hat{\mu}_{i}=\frac{1}{|S_{i}|}\cdot\sum_{p\in S_{i}}p$
  Output $\textsc{Aggregate}(\hat{\mu}_{1},\ldots,\hat{\mu}_{b\log\delta^{-1}})$
The sample complexity of this algorithm turns out to be optimal, as shown in Section 5. Phrased as a generalization question for bounding the excess risk given independent samples $S$ from an underlying arbitrary but fixed distribution, we show an optimal convergence rate of $\Theta\left(\frac{\log\delta^{-1}}{|S|}\right)$.
The remaining question is to find a good solution (i.e., run Aggregate) quickly. Clearly, the sample complexity is also a lower bound for the running time of any algorithm. Achieving this running time and in particular, breaking through the $O(\log^{2}\delta^{-1})$ barrier inherent in all previous algorithms, is a substantial challenge, even for computing a constant approximation, i.e. $\varepsilon\geq 1$. It is fairly easy to show that most of the empirical means are good estimators for the true mean. Selecting one of the good estimators is nevertheless difficult, as we have no way of estimating the cost and thus no easy way of distinguishing successful estimators from unsuccessful ones. Instead, we use one of the following two algorithms.
Geometric-Median-of-Means.
Our first algorithm is based on a generalization of the median-of-means estimator for the $1$-dimensional case. It turns out that in high dimensions, the corresponding analogue is to use the geometric median of $O(\log\delta^{-1})$ empirical means each of a sample of size $O(\varepsilon^{-1})$ yields a $(1+\varepsilon)$ approximate estimate for the mean, which is, in our view, a nice structural result (see Corollary 3.2 for a precise statement). Black box applications of gradient descent methods do not offer fast running times to compute suitable medians in our setting, but we show that a relatively simple gradient descent (see Algorithm 2) with a proper initialization finds a suitable solution in time $O((\varepsilon^{-1}+\log\log\delta^{-1})\cdot\log\delta^{-1}\cdot d)$.
MinSumSelect.
Our other algorithm, given in Algorithm 3, uses a carefully chosen order statistic combined with clustering ideas. Specifically, the observation is that the sum of distances of the $t$ closest points allows us to distinguish between a successful estimate and an unsuccessful estimate, for an appropriate choice of $t$. Here, the running time improves to $O((\varepsilon^{-1}+\log^{\gamma}\delta^{-1})\cdot\log\delta^{-1}\cdot d)$, for any constant $\gamma>0$. The algorithm can also be parameterized to achieve a similar running time to the gradient descent algorithm, at the cost of increasing the sample size by $\textmd{{poly}}\left(\log\log\delta^{-1}\right)$ factors.
1.2 Related Work
We covered the related work specific to sublinear algorithms for estimating the mean earlier, but there are several other closely related topics. Given a distribution, a classical question is to design good estimators for some underlying statistic. For high-dimensional means, there has been substantial attention recently, see (Lee & Valiant, 2022; Lugosi & Mendelson, 2019, 2024) and references therein. There are some common notions with our work, as the distance to the true mean and tail bounds for the estimators are an important quantity in our setting as well. Nevertheless, this line of work achieves stronger guarantees that typically require some underlying distributional assumption and the running time, which is a very important focus for sublinear algorithms, is not a primary concern.
In a more general setting, one could ask for approximate power means minimizing the objective $\min_{c\in\mathbb{R}^{d}}\sum_{p\in A}\|p-c\|^{z}$ for some positive integer $z$. For $z>2$, this problem was first posed by Cohen-Addad et al. (2021) who showed that $\tilde{O}(\varepsilon^{-z-3}\log^{2}\delta^{-1})$ samples were sufficient for a high success probability while $\Omega(\varepsilon^{-z+1})$ samples were necessary for constant success probability. Woodruff & Yasuda (2024) improved this even further, obtaining a sample complexity of $O(\varepsilon^{-z+1}(\log\varepsilon^{-1}+\log\delta^{-1})\log\delta^{-1})$. Perhaps surprisingly, the case $z=1$, also known as the geometric median requires $\Theta(\varepsilon^{-2}\log\delta^{-1})$ many samples, see (Chen & Derezinski, 2021; Parulekar et al., 2021) for lower bounds and Cohen et al. (2016) for an optimal algorithm. Finally, the special case $z\rightarrow\infty$ corresponds to the minimum enclosing ball problem. The solution is highly susceptible to outliers, thus any sublinear algorithm must either be allowed to discard a fraction of the input, or rely on additive guarantees (see (Clarkson et al., 2012) and (Ding, 2020) for examples).
[…]
2 Preliminaries
For a set of points $A\subset\mathbb{R}^{d}$, whose cardinality is $|A|=n$, we use $\mu(A):=\frac{1}{n}\cdot\sum_{p\in A}p$ to denote the mean, using $\mu$ if $A$ is clear from the context. Denote by $\textup{{Opt}}=\min_{c\in\mathbb{R}^{d}}\sum_{p\in A}\|p-c\|^{2}$. In addition, for a set of points $S$ sampled uniformly at random from $A$, we call $\hat{\mu}(S)$ the empirical estimator of $\mu(A)$, or simply $\hat{\mu}$ if $S$ is clear from context. A useful relationship between the goodness of $\hat{\mu}$ and the size of $|S|$ is established via the following identity.
[…]
$\mathbb{E}\left[\|\mu(A)-\hat{\mu}(S)\|^{2}\right]=\frac{1}{|S|}\cdot\frac{\textup{{Opt}}}{n}.$
At the heart of most algorithms for Euclidean means lies the following identity, which is arguably folklore.
[…]
For any set of points $A\subset\mathbb{R}^{d}$ and any $c\in\mathbb{R}^{d}$, we have
$\sum_{p\in A}\|p-c\|^{2}=\sum_{p\in A}\|p-\mu\|^{2}+n\cdot\|\mu-c\|^{2}.$
Lemma 2 implies that $\hat{\mu}$ is a $(1+\varepsilon)$ appoximation if and only if $\|\hat{\mu}-\mu\|\leq\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}$, which we will often use throughout this paper.
For the sake of completeness, we give a proof of the lemma.
Proof.
We have
$\sum_{p\in A}$ $\|p-c\|^{2}=\sum_{p\in A}\|p-\mu+\mu-c\|^{2}$
$=\sum_{p\in A}\left(\|p-\mu\|^{2}+\|\mu-c\|^{2}-2(p-\mu)^{\top}(\mu-c)\right)$
$=\sum_{p\in A}\|p-\mu\|^{2}+n\cdot\|\mu-c\|^{2},$
where the last equality follows from $\sum_{p\in A}(p-\mu)=0$. ∎
3 Geometric Median-of-Means and Gradient Descent
In this section, we show that the geometric median-of-means estimator has an optimal sample complexity and give a gradient descent algorithm (Algorithm 2) for computing a sufficiently good estimate efficiently. The gradient of the geometric median objective $\sum_{p\in P}\|p-c\|$ at any point $q$ with respect to some reference point set $P$ is $\nabla(q)=\sum_{p\in P}\nabla_{p}(q)$, where $\nabla_{p}(q):=\frac{q-p}{\|q-p\|}$ is the contribution to the gradient from point $p$ [^2] To make the gradient well defined, we exclude any points $p$ that are co-located with $q$.
Algorithm 2 FastGD$(P)$
  Compute coordinate-wise median $c_{0}$
  for $j=1,\ldots,T$ do
     Compute the gradient $\nabla(c_{j-1})=\sum_{p\in P}\frac{c_{j-1}-p}{\|c_{j-1}-p\|}$
     Project all $p_{i}\in P$ onto $c_{j-1}-\nabla(c_{j-1})$, denoting the projection by $p_{i,c_{j-1}}$
     Compute the $1$-dimensional median of the $p_{i,c_{j-1}}$’s along the line $c_{j-1}-\nabla(c_{j-1})$, denoting it by $c_{j}$
  Output $c_{T}$
The proof that the geometric median-of-means estimator is optimal is algorithmic, i.e. it will follow as a result of the analysis of Algorithm 2. We initialize Algorithm 2 with $P=\{\hat{\mu}_{1},\ldots,\hat{\mu}_{b\log\delta^{-1}}\}$ and $c_{0}$ as the coordinate-wise median, that is $(c_{0})_{i}$ is the median value of all points in $P$ along dimension $i$. In words, we compute $\nabla(q)$, and then execute a fast line search along $q-\nabla(q)$ by way of selecting the $1$-dimensional median of the projections of empirical means along $q-\nabla(q)$.
The goal is to prove the following theorem:
Theorem 3.1.
Algorithm 1 run with Algorithm 2 as an aggregation routine outputs a $(1+\varepsilon)$-approximate Euclidean mean with probability $1-\delta$ if $T\geq\gamma\cdot\log\log\delta^{-1}$ for some absolute constant $\gamma$. The running time is
$O\left(\left(\varepsilon^{-1}+\log\log\delta^{-1}\right)\cdot\log\delta^{-1}\cdot d\right).$
Proof Outline.
For the proof, we set the constants as follows - $a=1440$, $b=50$ and define $r:=\frac{1}{11}\cdot\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}$. The proof of Theorem 3.1 is articulated in three main steps. We first argue that, if $\|c_{j}-\mu\|\geq 10r$, then it must happen that, $\|c_{j+1}-\mu\|\leq 0.7\cdot\|c_{j}-\mu\|$, namely the statement of Lemma 3 (b). Because of a $O(\log\delta^{-1})^{2}$ approximation during initialization (see Lemma 3), after $T\in\Theta(\log\log\delta^{-1})$ iterations we converge to a point $c_{j}$ such that it is within a $10r$ radius around $\mu$. Once we have reached this point, we show, in Lemma 3, that successively chosen points could potentially escape such a ball by a wider ball of radius at most $11r$, but no further. In both cases, Algorithm 2 achieves a $(1+\varepsilon)$-approximation to the true Euclidean mean $\mu$.
We require a basic probabilistic lemma and a few common notions used in the remaining proof. We say that an empirical mean $\hat{\mu}_{i}$ is good, if $\|\hat{\mu}_{i}-\mu\|\leq r$. We denote the set of good empirical means by $G$, and define event
:=$\mathcal{E}:=\left\{|G|\geq\frac{7}{10}\cdot b\log\delta^{-1}\right\}.$ (1)
The following lemma is a straightforward application of the Chernoff bound.
[…]
With probability $1-\delta$, event $\mathcal{E}$ holds, for a sufficiently large absolute constants $a$ and $b$.
Proof.
Recall that the values of the constants were set as $a=1440$, $b=50$ and $r=\frac{1}{11}\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}$. Due to Lemma 2, Markov’s inequality and our choice of $a$, with probability at most $0.1$, we have $\|\hat{\mu}_{i}-\mu\|^{2}>\frac{\varepsilon\textup{{Opt}}}{121n}=r^{2}$, which implies that $\mathbb{E}[|G|]\geq\left(\frac{9}{10}\right)\cdot b\log\delta^{-1}$. For $b\geq 50$ we therefore get,
$\mathbb{P}\left[|G|\leq\frac{7}{10}\cdot b\log\delta^{-1}\right]\leq\text{exp}\left(-\frac{2}{81}\cdot b\log\delta^{-1}\right)\leq\delta,$
where the last inequality follows from a direct application of the Chernoff bound. ∎
Step 1: Fast convergence to close estimates from far ones.
In order to estimate the mean quickly with gradient descent, we will make use of the following geometric median gradient properties whenever our current estimate $c_{j}$ is far from $\mu$.
[…]
Conditioned on event $\mathcal{E}$, we have:
  1. (a)
    For any point $q\in\mathbb{R}^{d}$ with $\|q-\mu\|>10r$, then
    $\left\|\nabla(q)\right\|>\frac{3}{10}\cdot b\log\delta^{-1};$
  2. (b)
    If, at iteration $j$, Algorithm 2 chooses a point $c_{j}$ such that $\|c_{j}-\mu\|\geq 10r$, then
    $\|c_{j+1}-\mu\|\leq 0.7\cdot\|c_{j}-\mu\|.$
Proof.
We first show that every vector corresponding to a good mean is approximately in the same direction from $q$. Let $\hat{\mu}$ be a good solution. Consider the triangle $\{q,\mu,\hat{\mu}\}$ and let $\alpha_{i}$ and $\beta_{i}$ be the angle in $q$ and $\hat{\mu}$ respectively. We show an upper-bound for $cos(\alpha_{i})$. Using the law of sines,
$\frac{\sin(\alpha_{i})}{\left\|\mu-\hat{\mu}_{i}\right\|}=\frac{\sin(\beta_{i})}{\left\|q-\mu\right\|}\leq\frac{1}{\left\|q-\mu\right\|}$
We get $\sin(\alpha_{i})\leq\left\|\mu-\hat{\mu}_{i}\right\|/\left\|q-\mu\right\|\leq 0.1$, which implies $\cos(\alpha_{i})\geq\sqrt{(1-\sin^{2}(\alpha_{i})}\geq 0.99$.
To bound $\left\|\nabla(q)\right\|$, we consider the contribution of good and bad means separately. We write $\nabla(q)=\sum_{i\in[b\log\delta^{-1}]}\nabla_{i}(q)=\nabla_{G}(q)+\nabla_{B}(q)$, where $\nabla_{G},\nabla_{B}$ respectively denote good and bad points contributions to the gradient evaluated at $q$ (see Figure 1).
$\|\nabla_{G}(q)\|$ $=\left\|\sum_{i:~{}\hat{\mu}_{i}\text{ good}}\nabla_{i}(q)\right\|$
$\geq\left\|\sum_{i:~{}\hat{\mu}_{i}\text{ good}}\cos(\alpha_{i})\cdot\frac{\mu-q}{\|\mu-q\|}\right\|$
$\geq\frac{7}{10}\cdot 0.99\cdot b\log\delta^{-1},$
where the first inequality holds because projecting the gradient onto any direction only decreases the norm. Conversely, the contribution given by bad empirical means to the gradient norm is at most
$\|\nabla_{B}(q)\|$ $=\left\|\sum_{i:~{}\hat{\mu}_{i}\text{ not good}}\nabla_{i}(q)\right\|\leq\frac{3}{10}\cdot b\log\delta^{-1}$
where the inequality follows from the upper-bound on the number of bad means when $\mathcal{E}$ holds and the triangle inequality. Part (a) follows by combining both bounds: We have, due to the triangle inequality, an overall lower bound on the norm of the gradient of
$\left\|\sum_{i}\nabla_{i}(q)\right\|$ $\geq\left\|\sum_{i:~{}\hat{\mu}_{i}\text{ good}}\nabla_{i}(q)\right\|-\left\|\sum_{i:~{}\hat{\mu}_{i}\text{ not good}}\nabla_{i}(q)\right\|$
$\geq\left(\frac{7\cdot 0.99}{10}-\frac{3}{10}\right)\cdot b\log\delta^{-1}$
$\geq\frac{3}{10}\cdot b\log\delta^{-1}.$
For part (b), similarly to the above, we focus on the triangle $\{q,q-\nabla_{G},q-\nabla_{G}-\nabla_{B}\}$. Let $\beta$ be the angle in $q$ (see Figure 1, with $q$ instead of $c_{j}$) and let $\theta$ be the angle in $q-\nabla_{G}-\nabla_{B}$. Using the bounds of $\|\nabla_{G}\|$ and $\|\nabla_{B}\|$ and the law of sines, we get
$\frac{\sin\beta}{\|\nabla_{B}\|}=\frac{\sin\theta}{\|\nabla_{G}\|}\leq\frac{1}{\|\nabla_{G}\|}$
which implies $\sin\beta\leq{\|\nabla_{B}\|}/{\|\nabla_{G}\|}\leq 0.5$
Next, we upper-bound $\|c_{j+1}-\mu\|$ to prove part (b). Consider the projection of $\mu$ onto $c_{j}-\nabla_{G}(c_{j})$ (the green line in Figure 1), call it $y$. Now, let us consider the projection of $y$ onto $c_{j}-\nabla(c_{j})$, call it $\textmd{{proj}}_{j}(y)$. We know that $c_{j+1}$, the median of all the projected samples means (both good and bad), lies in the convex hull of good projected sample means. This implies that, for $c_{j+1}$, we must have
$\|c_{j+1}-\mu\|\leq\|c_{j+1}-y\|+\|y-\mu\|$
$\leq\underbrace{\|c_{j+1}-\textmd{{proj}}_{j}(y)\|}_{\leq r}+\underbrace{\|\textmd{{proj}}_{j}(y)-y\|}_{=\sin(\beta)\cdot\|c_{j}-y\|}+\underbrace{\|y-\mu\|}_{\leq r}$
$\leq\sin(\beta)\cdot\|c_{j}-\mu\|+2r$
$\leq 0.7\cdot\|c_{j}-\mu\|$
where the third inequality follows from the fact that $\|c_{j}-y\|\leq\|c_{j}-\mu\|$ (given that $c_{j}-\mu$ and $c_{j}-y$ are respectively the hypotenuse and the leg of the right-angle triangle $\{c_{j},\mu,y\}$ and $\|c_{j}-\mu\|\geq 10r$. ∎
Step 2: Once the estimate is close, it remains close.
[…]
Conditioned on event $\mathcal{E}$, if, at iteration $j$, Algorithm 2 chooses a point $c_{j}$ such that $\|c_{j}-\mu\|\leq 10r$, then
$\|c_{j+1}-\mu\|\leq 11r.$
Proof.
Consider a point $c_{j}$ chosen by Algorithm 2 such that $\|c_{j}-\mu\|\leq 10r$. Its gradient vector is $\nabla(c_{j})$ and Algorithm 2 projects all other empirical means onto $c_{j}-\nabla(c_{j})$ and takes the median to be $c_{j+1}$. For every mean $\hat{\mu}$, we denote its projection by proj$(\hat{\mu})$. We make two observations. Firstly, the set of proj$(\hat{\mu})$ of all good points lie on a bounded line segment of length at most $2r$. (see Figure 1). Second, due to the event $\mathcal{E}$, we have that the median chosen lies within the bounded line segment.
From the second observation, we have $\|c_{j+1}-\text{proj}(\mu)\|\leq r$, which is the radius of the good ball (see Figure 1).
Applying triangle inequality, we get
$\|c_{j+1}-\mu\|$ $\leq\|c_{j+1}-\text{proj}(\mu)\|+\|\text{proj}(\mu)-\mu\|$
$\leq r+\|c_{j}-\mu\|\leq 11r$
where the second inequality follows from the fact that $\text{proj}(\mu)$ is a projection on a line passing through $c_{j}$. ∎
$y$proj($y$)proj$(\mu)$$c_{j+1}$$(\hat{\mu}_{1})$$(\hat{\mu}_{2})$$c_{j}$$\mu$$-\nabla(c_{j})$$-\nabla_{G}(c_{j})$$-\nabla_{B}(c_{j})$$\beta$
Figure 1: Good empirical means are represented in green and bad ones in red. The ball of good means centered at $\mu$ has radius $r$. Projection of all the good means lie on the bounded line segment of length at most $2r$.
Step 3: Initialization.
[…]
Conditioned on event $\mathcal{E}$, let $m$ be the coordinate-wise median. Then
$\|m-\mu\|\leq b\cdot\log\delta^{-1}\cdot r$
Proof.
First, we note that each coordinate of $m$ lies in the convex hull of the projections of the good empirical means along that coordinate. This follows directly from the fact that $\mathcal{E}$ holds.
Second, each coordinate of $m$ is borrowed from a point, i.e., for every $i\in[d]$, there exists a mean $\hat{\mu}$ such that $m_{i}=\hat{\mu}_{i}$. We partition the coordinates into sets $S_{\hat{\mu}_{1},\hat{\mu}_{2}}$, where the $\{\hat{\mu}_{1},\hat{\mu}_{2}\}$ ranges over all pairs of good means. We say that $i\in S_{\hat{\mu}_{1},\hat{\mu}_{2}}$ if $\hat{\mu}_{1_{i}}\leq m_{i}\leq\hat{\mu}_{2_{i}}$ or $\hat{\mu}_{2_{i}}\leq m_{i}\leq\hat{\mu}_{1_{i}}$. Looking at a fixed pair, we get
$\sum_{i\in S_{\hat{\mu}_{1},\hat{\mu}_{2}}}(m_{i}-\mu_{i})^{2}$ $\leq\sum_{i\in[d]}(m_{i}-\hat{\mu}_{1_{i}})^{2}+(m_{i}-\hat{\mu}_{2_{i}})^{2}$
$\leq 2\cdot r^{2}$
Summing the inequality up over all unordered pairs $\{\hat{\mu}_{1},\hat{\mu}_{2}\}$ in $G$ (there are at most $1/2\cdot(b\cdot\log\delta^{-1})^{2}$ pairs), we get
$\|m-\mu\|^{2}$$\leq\sum_{\{\hat{\mu}_{1},\hat{\mu}_{2}\}}\sum_{i\in S_{\hat{\mu}_{1},\hat{\mu}_{2}}}(m_{i}-\mu_{i})^{2}$
$\leq(b\cdot\log\delta^{-1})^{2}\cdot r^{2}$
as desired. ∎
We are now ready to prove our main result.
Proof of Theorem 3.1.
We condition on event $\mathcal{E}$. In Algorithm 2, the initial value $c_{0}$ is taken to be the coordinate-wise median $m$ of all the $b\log\delta^{-1}$ sample means. From Lemma 3, we know that $\|m-\mu\|\leq\left(b\cdot\log\delta^{-1}\right)\cdot r$. Using part (b) of Lemma 3, we have that as long as $\|c_{j}-\mu\|\geq 10r$, it holds that $\|c_{j+1}-\mu\|\leq 0.7\cdot\|c_{j}-\mu\|$. As the quantity $\|c_{j}-\mu\|$ is decreasing exponentially, it follows that for some absolute constants $\gamma_{1},\gamma_{2}$ there exists a $j\leq\gamma_{1}+\gamma_{2}\log\log\delta^{-1}$, such that $\|c_{j}-\mu\|\leq 10r$. In the next iterations of the algorithm, from Lemma 3, we know that for the terminating value of the algorithm, $c_{T}$, we have $\|c_{T}-\mu\|\leq 11r$, which yields the desired approximation ratio.
Concerning the the running time, we first need to compute $b\log\delta^{-1}$ empirical means from $a\varepsilon^{-1}$ many uniform samples, and then, in each of the $T$ iterations, there are $b\log\delta^{-1}$ sample means to project onto the gradient vector, an operation that takes $O(d\cdot\log\delta^{-1})$ time, with $d$ being the dimension of the ambient space. To compute the median of all projections in every iteration, as well as finding the initial coordinate-wise median $c_{0}$ can be done using a linear-time rank select procedure (see (Cormen et al., 2009, Chapter 9.3)). For each iteration, as well as the initialization, this therefore takes $O(\log\delta^{-1}\cdot d)$ time. Finally, using $T=O(\log\log\delta^{-1})$, the overall running time is
$O\left(\left(\varepsilon^{-1}+\log\log\delta^{-1}\right)\cdot\log\delta^{-1}\cdot d\right),$
which concludes the proof. ∎
Optimality of the Geometric Median-of-Means Estimator.
Finally, we wish to briefly remark on the sample complexity of the geometric median-of-means estimator and generalization bounds. Optimality follows from part (a) of Lemma 3, as if the gradient of the geometric median $m$ is 00, $\|m-\mu\|\leq 10r$ must hold.
Corollary 3.2.
For sufficiently large absolute constants $a$ and $b$, the geometric median $m=\underset{c\in\mathbb{R}^{d}}{\arg\min}\sum_{i=1}^{b\log\delta^{-1}}\|\hat{\mu}_{i}-c\|$ of $b\log\delta^{-1}$ many empirical means consisting of $a\varepsilon^{-1}$ many points is a $(1+\varepsilon)$-approximate mean with probability at least $1-\delta$.
Remark 3.3.
The geometric median cannot be computed exactly, we can merely approximate it. The quality of the choice of approximation is important here. Using event $\mathcal{E}$ by itself, it is not clear that even a $2$-approximate geometric median will recover a $(1+O(\varepsilon))$-approximate mean. Indeed, if we assume that a single $\hat{\mu}_{i}$ has distance $\sqrt{\frac{\varepsilon\cdot\textup{{Opt}}}{\delta n}}$ to the optimal mean, we must compute a $(1+\delta^{-O(1)})$-approximate median. In Section 5, we will show that this assumption is warranted for worst case instances. The fastest algorithm for finding a $(1+\gamma)$-approximate geometric median by (Cohen et al., 2016) runs in time $O(nd\log^{3}\gamma^{-1})$. In our setting, $n=b\log\delta^{-1}$. The resulting running time of $O(\log^{4}\delta^{-1}\cdot d)$ is thus substantially slower that the other two algorithms presented in this paper.
Generalization Bounds.
[…]
Corollary 3.4.
Given a set of independent samples $S$ drawn from some underlying arbitrary but fixed distribution supported on $B_{2}^{d}$, the geometric median-of-means estimator has an excess risk for the least squared distances objective of
^$\hat{\mathcal{R}}-\mathcal{R}\leq\gamma\cdot\frac{\log\delta^{-1}}{|S|}$
with probability $1-\delta$ for some absolute constant $\gamma>0$.
[…]
4 Mean Estimation via Order Statistics
In this section, we leverage order statistics of the candidate means so as to allow for a quicker aggregation of a suitable candidate mean. The main result is as follows.
Algorithm 3 $\textsc{MinSumSelect}(P,i)$
  Input: Set of points $p_{1},\ldots p_{|P|}$, recursion depth $i$
  if $i=0\textbf{ or }|P|=1$  then
     $W\leftarrow P$
  else
     Split $P$ arbitrarily into $\sqrt{|P|}$-sized clusters $\{P_{1},\ldots,P_{\sqrt{|P|}}\}$
     $W\leftarrow\emptyset$
     for each $P_{j}$ do
        $W\leftarrow W\cup\{\textsc{MinSumSelect}(P_{j},i-1)\}$
  Output $\textsc{ComputeWinner}(W)$
Algorithm 4 ComputeWinner($P$)
  for each $p_{j}\in P$ do
     Let $p_{j}^{\prime}\in P$ be the $\frac{7}{10}|P|$-closest point to $p_{j}$
     Compute $D_{j}:=\sum_{p\in P,\|p-p_{j}\|\leq\|p_{j}^{\prime}-p_{j}\|}\|p-p_{j}\|$
  Output $\arg\min_{p_{j}\in P}D_{j}$
Theorem 4.1.
Algorithm 1 run with Algorithm 3 as an aggregation routine outputs a $(1+\varepsilon)$ approximation with probability $1-\delta$, using a sample of size $O\left(100^{i}\varepsilon^{-1}\log\delta^{-1}\right)$, and running in time
$O\left(\left(100^{i}\varepsilon^{-1}+\log^{2^{-i}}\delta^{-1}\right)\log\delta^{-1}\cdot d\right),$
for any non-negative integer $i$.
The key observation is that a good mean can always be identified via ComputeWinner as an estimate $\hat{\mu}_{j}$ minimizing the sum of distances of the $\frac{7}{10}$th means closest to $\hat{\mu}_{j}$. In a nutshell, any mean minimizing such a sum must be close to sufficiently many successful estimates. Unfortunately, a naive implementation of this idea takes time $O(\log^{2}\delta^{-1}\cdot d)$, as proven in the following lemma.
[…]
ComputeWinner($P$) takes time $O(|P|^{2}\cdot d)$.
Proof.
We compute all pairwise distances between the points in $P$, which takes time $O(|P|^{2}\cdot d)$. To compute $D_{j}$, we first have to find the $\frac{7}{10}|P|$-closest point to $p_{j}$, which takes time $O(|P|)$ with a sufficiently good rank select procedure, see (Cormen et al., 2009, Chapter 9.3). Thereafter, summing up all the distances takes time $O(|P|)$ per $p_{j}\in P$. ∎
Nevertheless, these scores yield an improved running time by arbitrarily partitioning the estimates into $\sqrt{|P|}$ groups, finding a good estimate in every group via the truncated sum statistic each in time $|P|\cdot d$, for a total running time of $|P|^{3/2}\cdot d$, and then selecting the best estimate via another truncated sum statistic on all estimates returned by the groups in time $|P|$. The final algorithm consists of applying this idea recursively.
To prove that MinSumSelect outputs a good solution, we require a parameterized notion of a successful empirical mean. We say that a mean $\mu_{j}$ is $\gamma$-good, if $\|\mu_{j}-\mu\|\leq\gamma\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}$. A straightforward application of the Chernoff bound guarantees us that all but a very small fraction of the points are $\gamma$-good.
The following lemma determines the quality of the computed solution.
[…]
Let $P$ be a set of means such that at least $\left(1-\left(\frac{3}{10}\right)^{i+1}\right)$ of the means are $\gamma$-good. Then $\textup{{MinSumSelect}}(P,i)$ returns a mean that is $5^{i+1}\gamma$-good.
We prove this lemma by induction. We will use the following lemma in both the base case and the inductive step.
[…]
Given a set of means $P$, suppose that at least $\frac{7}{10}|P|$ are $\gamma$-good. Then, the estimate returned by ComputeWinner($P$) is $5\gamma$-good.
Proof.
First, let $\hat{\mu}^{\prime}\in P$ be any $\gamma$-good estimator. We know that the $\frac{7}{10}|P|$-closest estimators to $\hat{\mu}^{\prime}$ have distance at most $2\gamma\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}$, which likewise implies that $\min_{\mu_{j}\in P}D_{j}\leq\frac{7}{10}|P|\cdot 2\gamma\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}$. Now suppose that $\hat{\mu}_{j}$ is the estimator returned by the algorithm. Let $G(\hat{\mu}_{j})$ be the set of $\gamma$-good estimators among the $\frac{7}{10}|P|$ closest to $\hat{\mu}_{j}$. By assumption, we have $|G(\hat{\mu}_{j})|\geq\frac{4}{10}|P|$, which implies
$\sum_{\hat{\mu}_{k}\in G(\hat{\mu}_{j})}\|\hat{\mu}_{j}-\hat{\mu}_{k}\|$ $\leq\frac{7}{10}|P|\cdot 2\gamma\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}$
$=\frac{7}{5}\gamma|P|\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}},$
which gives
$\min_{\hat{\mu}_{k}\in G(\hat{\mu}_{j})}\|\hat{\mu}_{j}-\hat{\mu}_{k}\|\leq\frac{7}{2}\gamma\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}.$
Therefore,
$\|\hat{\mu}_{j}-\mu\|$ $\leq\min_{\hat{\mu}_{k}\in\textup{succ}(\hat{\mu}_{j})}\|\hat{\mu}_{j}-\hat{\mu}_{k}\|+\|\hat{\mu}_{k}-\mu\|$
$\leq\frac{7}{2}\gamma\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}+\gamma\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}}\leq 5\gamma\sqrt{\frac{\varepsilon\textup{{Opt}}}{n}},$
as desired. ∎
Proof of Lemma 4.
We proceed with the induction starting from $i=0$.
Base Case. For the base case $i=0$, MinSumSelect($P$,00) only calls ComputeWinner. Thus, this case holds due to Lemma 4.
Inductive Step. Let $P_{1},\ldots P_{\sqrt{|P|}}$ be the clusters of $P$ computed when first calling $\textsc{MinSumSelect}(P,i)$. By assumption, we have at most $\left(\frac{3}{10}\right)^{i+1}|P|$ means that are not $\gamma$-good. This implies that the number of clusters with more than $\left(\frac{3}{10}\right)^{i}\sqrt{|P|}$ means that are not $\gamma$-good is at most
$\frac{\left(\frac{3}{10}\right)^{i+1}|P|}{\left(\frac{3}{10}\right)^{i}\sqrt{|P|}}=\frac{3}{10}\sqrt{|P|}.$
Denote this set by $B(P)$ and let $G(P)$ be the remaining clusters. For each $P_{j}\in B(P)\cup G(P)$, let $\hat{\mu}_{j}$ returned by MinSumSelect($P_{j}$,$i-1$). If $P_{j}\in G(P)$, we may use the inductive hypothesis which states that the mean $\hat{\mu}_{j}$ returned by MinSumSelect($P_{j}$,$i-1$) is $5^{i}\gamma$-good. Since at least $\frac{7}{10}\sqrt{|P|}$ of the thus computed means are $5^{i}\gamma$-good, we may use Lemma 4 which shows that the final mean returned by $\textsc{ComputeWinner}(\cup_{P_{j}\in B(P)\cup G(P)}\{\hat{\mu}_{j}\})$ is $5^{i+1}\gamma$-good, which concludes the proof. ∎
To conclude the proof, we require two more arguments. First, we must show that, for an appropriate choice of $a$ and $b$, at least a $\left(1-\left(\frac{3}{10}\right)^{i+1}\right)$ fraction of the means are $1$-good with probability at least $1-\delta$, allowing us to use Lemma 4. Second, we will argue the running time. The first is a simple application of the Chernoff bound.
[…]
For $a\geq 2\cdot 25^{i+1}\cdot\left(\frac{10}{3}\right)^{i+1}$ and $b=3$, at least $\left(1-\left(\frac{3}{10}\right)^{i+1}\right)$ of the means are $5^{-(i+1)}$-good with probability $1-\delta$.
Proof.
Let $X_{i}=\begin{cases}1&\text{if }\mu_{i}\text{ is not }5^{-(i+1)}-\text{good}\\ 0&\text{if }\mu_{i}\text{ is }5^{-(i+1)}-\text{good}\end{cases}$.
We have $\mathbb{E}[\|\hat{\mu}_{i}-\mu\|^{2}]=\frac{1}{a}\cdot\frac{\varepsilon\textup{{Opt}}}{n}$ due to Lemma 2. This implies due to Markov’s inequality that $\mathbb{P}[X_{i}=1]\leq\frac{25^{i+1}}{a}$ which, by our choice of $a$, is less than $\frac{1}{2}\cdot\left(\frac{3}{10}\right)^{i+1}$. Thus, by the Chernoff bound
$\mathbb{P}\left[\sum_{i=1}^{b\log\delta^{-1}}X_{i}\geq\left(\frac{3}{10}\right)^{i+1}b\log\delta^{-1}\right]$
$\leq\mathbb{P}\left[\sum_{i=1}^{b\log\delta^{-1}}X_{i}\geq 2\cdot\frac{25^{i+1}}{a}\cdot b\log\delta^{-1}\right]$
$\leq{\textup{exp}}\left(-\frac{25^{i+1}}{3a}\cdot b\log\delta^{-1}\right),$
which is at most $\delta$ by our choice of $b$. ∎
We now give a complete proof of Theorem 4.1.
Proof of Theorem 4.1.
Throughout this proof, assume $a\geq 2\cdot 25^{i+1}\left(\frac{10}{3}\right)^{i+1}$ and $b\geq 3$.
We first argue correctness, then running time. Let $P$ denote the entire set of means passed to the MinSumSelect algorithm. By Lemma 4 and Lemma 4 and with our choices of $a$ and $b$, $\textsc{MinSumSelect}(P,i)$ returns a $1$-good mean, that is a $(1+\varepsilon)$-approximate mean with probability at least $1-\delta$.
What remains to be shown is the running time. Denote by $|P|_{0}=b\log\delta^{-1}$ the initial set of empirical means. $\textsc{MinSumSelect}(P,i)$ computes $\sqrt{|P|}$ children, each of which recursively calls MinSumSelect. Consolidating via ComputeWinner takes time $O(\sqrt{|P|}^{2}\cdot d)=O(|P|\cdot d)$ due to Lemma 4. Thus the overall recursion takes time
$T(|P|)=\begin{cases}\sqrt{|P|}\cdot T(\sqrt{|P|})+|P|\cdot d&\text{if }|P|\geq|P|_{0}^{2^{-i}}\\ |P|^{2}\cdot d&\text{if }|P|\leq|P|_{0}^{2^{-i}}\end{cases}$
which solves to a running time of $O(i\cdot|P|_{0}\cdot d+|P|_{0}^{1+2^{-i}}\cdot d)=O((b\log\delta^{-1})^{1+2^{-i}}\cdot d)$. The time to compute the initial set of means is $O\left(a\varepsilon^{-1}\cdot b\log\delta^{-1}\cdot d\right)$, which with our choice of $a$ and $b$, and some overestimation of the constants, leads to an overall running time of $O\left(\left(100^{i}\varepsilon^{-1}+\left(\log\delta^{-1}\right)^{2^{-i}}\right)\cdot\log\delta^{-1}\cdot d\right)$. ∎
Remark 4.2.
For any constant choice of recursion depth, the sample complexity only increases by constants. We did not attempt to optimize the constants, but the exponential dependency on $i$ is not avoidable. If we were to prioritize the running time over the sample complexity, we can set a recursion depth of $i=\log\log\log\delta^{-1}$, which achieves a running time and sample complexity of $O(\varepsilon^{-1}\log\delta^{-1}\textmd{{poly}}\left(\log\log\delta^{-1}\right)\cdot d)$.
Furthermore, if the recursion depth is shallow (i.e., for small values of $i$), the recursion can be improved via a better choice of cluster size. Specifically, in the case $i=1$, that is with just one set of children, Theorem 4.1 yields a running time of $O\left((\varepsilon^{-1}+\sqrt{\log\delta^{-1}})\cdot\log\delta^{-1}\cdot d\right)$. If we instead choose $\left(\log\delta^{-1}\right)^{2/3}$ many clusters, each consisting of $3]{\log\delta^{-1}}$ many estimators, we obtain a running time of $O\left((\varepsilon^{-1}+\sqrt[3]{\log\delta^{-1}})\cdot\log\delta^{-1}\cdot d\right)$. Similar improvements for other values $i$ are also possible, but these improvements become increasingly irrelevant compared to the bounds given in Theorem 4.1 as $i$ gets larger.
5 Lower Bounds
We complement our algorithmic results with lower bounds.
5.1 A High Probability Lower Bound for Mean-Estimation
Theorem 5.1.
Any sublinear algorithm that outputs a $(1+\varepsilon)$-approximate Euclidean mean with probability at least $1-\delta$ must sample at least $\Omega(\varepsilon^{-1}\log\delta^{-1})$ many points.
The idea is to generate two instances that a sublinear algorithm for approximating means has to distinguish between and then bound the number of samples required to distinguish between such distributions. The first instance is virtually identical to the one used by (Cohen-Addad et al., 2021). It places $n$ points at 00 and $\varepsilon n$ points at $1$. Thus, the optimal mean is $\frac{\varepsilon}{1+\varepsilon}$. A routine calculation via Lemma 2 shows that the optimal cost is $\frac{\varepsilon n}{1+\varepsilon}$. The second instance places all points at 00. Since 00 is not a sufficiently good approximate mean for the first instance, a sublinear algorithm can distinguish between the two instances, which yields a lower bound on the sample complexity.
Proof.
We give two instances. The first instance places $n$ points at 00 and $\varepsilon n$ points at $1$. Thus, the optimal mean is placed at $\frac{\varepsilon}{1+\varepsilon}$. A routine calculation via Lemma 2 shows that the optimal cost is $\textup{{Opt}}=\frac{\varepsilon n}{1+\varepsilon}$. The second instance places all points at 00.
First we argue that a sublinear algorithm can distinguish between these two instances. Any approximate mean for the second instance must output 00. If we output 00 for the first instance, we incur a cost of $\varepsilon n=\frac{\varepsilon n(1+\varepsilon)}{1+\varepsilon}=\textup{{Opt}}+\varepsilon\cdot\textup{{Opt}}$, hence any algorithm improving over a $(1+\varepsilon)$ approximation for the former cannot output 00. Thus, a necessary condition to distinguish between the two instances is for the algorithm to sample at least one point at $1$.
Suppose we sample $m$ points. Our goal is to show that for $m\in\Omega(\varepsilon^{-1}\log\delta^{-1})$, the probability that all of the sampled points are drawn from 00 is less than $\delta$ for the first instance. Let $X_{i}$ denote the indicator variable of the $i$th sampled point. We have
$\delta$ $\geq\mathbb{P}[\forall~{}i,~{}X_{i}=0]=\mathbb{P}[X_{1}=0]^{m}=\left(\frac{1}{1+\varepsilon}\right)^{m}$
$={\textup{exp}}\left(m\ln\frac{1}{1+\varepsilon}\right)={\textup{exp}}\left(-m\ln(1+\varepsilon)\right)$
$>{\textup{exp}}\left(-m\varepsilon\right),$
where the final inequality follows from the Mercator series $\ln(1+\varepsilon)=\sum_{i=1}^{\infty}\left(\frac{\varepsilon^{i}}{i}\right)\cdot(-1)^{i+1}$. Thus, for any $m<\varepsilon^{-1}\cdot\log\delta^{-1}$, we do not achieve the desired failure probability. Conversely, $m\in\Omega(\varepsilon^{-1}\log\delta^{-1})$ for an algorithm to succeed. ∎
The two instances lie within the unit Euclidean ball. By normalizing the number of points, it also yields a distribution supported on the unit Euclidean ball, which implies that the generalization bounds given in Corollary 3.4 are sharp.
5.2 The Empirical Mean is not a Good Estimator
Finally, we briefly show that the arguably most natural algorithm that outputs the empirical mean of a subsampled point set has a substantial increase in sample complexity in the high success probability regime and therefore also a substantial increase in running time compared to the algorithms presented in this paper. The result is likely folklore, but included for completeness.
Theorem 5.2.
For all $\varepsilon,\delta\in(0,1)$, there exists an instance such that $\Omega(\varepsilon^{-1}\delta^{-1})$ independently sampled points are required for the empirical mean to be a $(1+\varepsilon)$-approximate Euclidean mean with probability at least $1-\delta$.
The proof is based on reducing the problem of finding a good approximate mean with high probability to giving an example distribution for which the Chebychev inequality is tight.
Proof.
Consider a sample $|S|$, with $|S|$ being specified later. We generate an instance as follows. We place a $\frac{1}{2|S|^{2}\varepsilon}$-fraction of the points each at $-|S|\sqrt{\varepsilon}$ and $|S|\sqrt{\varepsilon}$ and the remaining $1-\frac{1}{|S|^{2}\varepsilon}$ fraction at 00. Then the optimal solution places the mean at 00 and has average cost $(|S|\sqrt{\varepsilon})^{2}\cdot\frac{1}{|S|^{2}\varepsilon}=1$. By Lemma 2, this implies that the empirical mean $\hat{\mu}=\frac{1}{|S|}\sum_{p\in S}p$ is a better than $(1+\varepsilon)$-approximate mean if and only if $|\hat{\mu}|<\sqrt{\varepsilon}$. We set the failure probability, that is the probability $\mathbb{P}[|\hat{\mu}|\geq\sqrt{\varepsilon}]=\delta$. Then $\mathbb{P}[|\hat{\mu}|\geq\sqrt{\varepsilon}]$ is at least the probability that we sample exactly one point from either $-|S|\sqrt{\varepsilon}$ or $|S|\sqrt{\varepsilon}$ and the remaining $|S|-1$ points from 00. Using Bernoulli’s inequality and the density function of the binomial distribution, we therefore have
$\delta=\mathbb{P}[|\hat{\mu}|\geq\sqrt{\varepsilon}]$ $\geq|S|\cdot\frac{1}{|S|^{2}\varepsilon}\cdot\left(1-\frac{1}{|S|^{2}\varepsilon}\right)^{|S|-1}$
$\geq\frac{1}{|S|\varepsilon}\cdot\left(1-\frac{1}{|S|\varepsilon}\right)$
Solving $\frac{1}{|S|\varepsilon}\cdot\left(1-\frac{1}{|S|\varepsilon}\right)\leq\delta$ for $|S|$ then yields $|S|\in\Omega(\varepsilon^{-1}\delta^{-1})$, as desired. ∎
Acknowledgements
Matteo Russo is partially supported by the ERC Advanced Grant 788893 AMDROMA “Algorithmic and Mechanism Design Research in Online Markets”. Beatrice Bertolotti and Chris Schwiegelshohn are partially supported by the Independent Research Fund Denmark (DFF) under a Sapere Aude Research Leader grant No 1051-00106B. Sudarshan Shyam is partially supported by the Independent Research Fund Denmark (DFF) under grant 2032-00185B.